OpenStudy (anonymous):

Solve for x. x - 7/8 = 7/2 A. x = 31 1/2 B. x = 35 C. x = 63 D. x = 9 2/7

OpenStudy (harsimran_hs4):

x = 7/2 + 7/8 x = (56 + 14)/16 x = 70/16 = 35/8 and it doesn`t match any so can you please use equation generator so that the question is clear?

OpenStudy (anonymous):

the what?

OpenStudy (harsimran_hs4):

make use of the brackets atleast to clarify the question

OpenStudy (anonymous):

He's right, they're all wrong. x=4.375 4 3/8

OpenStudy (anonymous):